# Stacked Intelligent Metasurfaces for Task-Oriented Semantic Communications

### v1

@@ Abstract
Semantic communication leveraging advanced deep learning (DL) technologies enhances the efficiency, reliability, and security of information transmission. Emerging stacked intelligent metasurface (SIM) having a diffractive neural network (DNN) architecture allows performing complex calculations at the speed of light. In this letter, we introduce an innovative SIM-aided semantic communication system for image recognition tasks. In the considered model, a SIM is positioned in front of the transmitting antenna. In contrast to conventional communication systems transmitting the modulated signals carrying the image information or compressed semantic information, the carrier electromagnetic (EM) wave is directly transmitted from the source in the proposed system. The input layer of the SIM is utilized for source encoding, while the remaining multi-layer architecture constitutes a DNN for semantic encoding. Specifically, the semantic encoder aims to transform the signals passing through the input layer of the SIM into a unique beam towards a receiving antenna corresponding to the image class. Remarkably, both the source and semantic encoding occur naturally as the EM waves propagate through the SIM. At the receiver, the image is recognized by probing the received signal magnitude across the receiving array. To this end, we develop an efficient algorithm to train the transmission coefficients of SIM’s meta-atoms to learn the semantic representation of the image. Extensive numerical results verify the effectiveness of utilizing the SIM-based DNN for image recognition task-oriented semantic communications, achieving more than 90% recognition accuracy.
@@
Semantic communication, stacked intelligent metasurface (SIM), image recognition, wave-domain computation.
@@ I Introduction
Conventional content-centric communication paradigms fail to satisfy the stringent quality-of-service (QoS) requirements of next-generation wireless networks, particularly in terms of ultra-reliability, low latency, and high data rates [1]. As a result, there is a growing interest in intent-driven networks, which triggers the recent research surge in semantic communications [2, 3]. Relying on advanced artificial intelligence (AI) technologies and powerful computing capability, semantic communication enables learning the semantic representations of various types of sources. For instance, the authors of [4] proposed a novel text semantic communication system, DeepSC, to recover the meaning of sentences at the receiver with the minimum semantic errors. Based on this, several variants of DeepSC and advanced deep learning (DL)-driven semantic communication systems have been developed [5, 6]. Nevertheless, the joint optimization of the semantic and channel encoders in these semantic communication systems is typically time-consuming, which undoubtedly increases the training costs and deployment challenges for edges [6]. Additionally, the inference process would suffer from high computational complexity as the number of nodes increases.
@@
Recently, stacked intelligent metasurface (SIM) emerged as a promising technique for realizing fast and efficient wave-domain computation. Specifically, a SIM is a closed vacuum container with multiple closely stacked metasurface layers, where each metasurface consists of numerous low-cost meta-atoms whose electromagnetic (EM) response can be flexibly tuned by an intelligent controller, e.g., a field programmable gate array (FPGA) [7]. As a result, SIM forms a programmable diffractive neural network (DNN) in the EM domain, with signal processing implemented at the speed of light [8, 9]. Recently, researchers have utilized SIMs to perform various computing tasks in wireless communication and sensing scenarios [10, 11, 12]. Specifically, the authors of [10] proposed a deep reinforcement learning (DRL) approach to optimize the transmission coefficients of the SIM’s meta-atoms to perform downlink precoding in the wave domain. Additionally, the authors of [13] utilized SIM to perform a two-dimensional discrete Fourier transform (DFT) to enable efficient direction-of-arrival (DOA) estimation. Nevertheless, the potential of utilizing SIM for semantic communications has not hitherto been investigated.
@@
Against this background, we examine an image recognition task-oriented semantic communication system leveraging a SIM. Compared to conventional semantic communication systems sequentially encoding and transmitting semantic information, the wave-based computing paradigm enables an automatic and tractable source and semantic encoding as the EM waves propagate through the SIM, which greatly reduces the energy consumption and processing latency. Specifically, in the proposed system, a SIM is positioned in front of the transmitting antenna, which only transmits the carrier EM waves. Furthermore, the input layer of the SIM is utilized for image source encoding, while the remaining multi-layer architecture constitutes a DNN for image semantic encoding. The semantic encoder aims to transform the signals passing through the input layer of the SIM into a unique beam towards an antenna corresponding to the image class. As a result, the receiver recognizes the image by probing the signal magnitude across the receiving array. To this end, we develop an efficient mini-batch gradient descent algorithm to train the transmission coefficients of meta-atoms in the SIM. Finally, extensive simulation results verify the effectiveness of utilizing a SIM to realize the image recognition task-oriented semantic communications.
@@
Figure 1: A SIM-aided image recognition task-oriented semantic communication system, where we consider $M$ classes.
@@ II System Model
In this section, we introduce the SIM-aided semantic communication system for image recognition. As shown in Fig. 1, we consider a system consisting of a single-antenna transmitter and a receiver equipped with $M$ antennas, which correspond to the number of image classes. Let =$\mathcal{M}=\left\{1,2,\ldots,M\right\}$ denote the set of indexes of antennas. More specifically, the transmitting antenna emits the unmodulated radio wave, and the SIM performs source and semantic coding of the image to direct the transmissive EM waves towards the corresponding antenna at the receiver.
@@ II-A1 Transmitting Antenna
In contrast to conventional communication systems, the transmitting antenna only radiates the carrier EM waves. Let $\sqrt{p_{t}}s$ denote the transmit signal, where $p_{t}$ represents the transmit power, and $s$ is the normalized EM signal, satisfying $|s|^{2}=1$.
@@ II-A2 SIM
$\begin{split}w^{l}_{n,n^{*}}&=f_{w}\left(d_{s},d_{n,n^{*}}^{l},S_{a},\lambda\right)\\ &=\frac{d_{s}S_{a}}{\left({d_{n,n^{*}}^{l}}\right)^{2}}\left(\frac{1}{2\pi d_{n,n^{*}}^{l}}-j\frac{1}{\lambda}\right)e^{j2\pi d_{n,n^{*}}^{l}},\end{split}$ (1)
for $l\in\mathcal{L}/\{0\}$, $n^{*},n\in\mathcal{N}$, where $S_{a}$ is the area of each meta-atom, $d^{l}_{n,n^{*}}$ denotes the corresponding propagation distance between the two meta-atoms, $\lambda$ denotes the radio wavelength. Moreover, since the SIM is closely placed in front of the transmitter, the propagation channel from the transmitting antenna to the input layer of the SIM is characterized by $\mathbf{w}^{0}=f_{w}\left(d_{t,s},d^{0}_{n},S_{a},\lambda\right)\in\mathbb{C}^{N\times 1}$ [9], where $d_{t,s}$ and $d^{0}_{n}$ denote the horizontal distance from the transmitting antenna to the input layer of the SIM and the distance from the transmitter to the $n$-th meta-atom on the input layer, respectively.
@@
As illustrated in Fig. 1, the input layer of the SIM serves as a digital-to-analog converter to transform the pixel matrix of the source image into the transmission coefficients of the meta-atoms on the input layer [^1] For the sake of brevity, we consider only greyscale images in this letter. For color images, each pixel matrix of three channels can be preprocessed for fully leveraging the amplitude and phase components of the transmission coefficients of meta-atoms on the input layer.. Specifically, the image is first preprocessed to match the size of the input layer $\left(N_{\mathrm{row}}\times N_{\mathrm{col}}\right)$ by employing appropriate extraction or interpolation techniques [16], and then the transmission coefficients of the source encoding layer $\left(z^{0}_{1},z^{0}_{2},\ldots,z^{0}_{N}\right)$ are configured to the corresponding value of the pixel matrix. The amplitude and phase of the EM waves passing through the source encoding layer are determined by the product of the incident signal on the source coding layer, i.e., $\sqrt{p_{t}}s\mathbf{w}^{0}$ and the EM response imposed by the input layer, i.e., $\mathbf{Z}^{0}$. Therefore, the source encoded signal $\mathbf{x}\in\mathbb{C}^{N\times 1}$ radiated from the input metasurface layer can be expressed as
@@
The rest layers of the SIM construct a DNN, which serves as a semantic encoder, aiming to transform $\mathbf{x}$ to a unique beam towards the receiving antenna corresponding to its image class. According to the Huygens–Fresnel principle [14], the EM wave passing through each meta-atom on the source encoding layer acts as a secondary source to illuminate all the meta-atoms on the first metasurface layer, and this process is repeated as the waves propagate through all metasurface layers [15]. Let $\mathbf{B}=\mathbf{Z}^{L}\mathbf{W}^{L}\ldots\mathbf{Z}^{2}\mathbf{W}^{2}\mathbf{Z}^{1}\mathbf{W}^{1}\in\mathbb{C}^{N\times N}$ denote the EM response of the semantic encoding layers. Hence, the semantic encoded signal $\tilde{\mathbf{x}}\in\mathbb{C}^{N\times 1}$ can be represented as
$\begin{split}\tilde{\mathbf{x}}=\mathbf{B}\mathbf{x}=\sqrt{p_{t}}s\mathbf{Z}^{L}\mathbf{W}^{L}\ldots\mathbf{Z}^{2}\mathbf{W}^{2}\mathbf{Z}^{1}\mathbf{W}^{1}\mathbf{Z}^{0}\mathbf{w}^{0}.\end{split}$ (3)
@@ II-B Channel Model
where $p$ is the path loss, $K$ is the Rician factor, $\mathbf{H}_{\mathrm{LOS}}$ is the line-of-sight (LOS) component, characterizing the direct path between SIM and receiver, which is defined by ${[\mathbf{H}_{\mathrm{LOS}}]}_{m,n}=e^{-j\frac{2\pi}{\lambda}d_{m,n}}$, where $d_{m,n}$ denotes the distance from the $n$-th meta-atom on the output layer of the SIM to the $m$-th antenna at the receiving array. $\mathbf{H}_{\mathrm{NLOS}}$ is the non-LOS (NLOS) component, following the circularly symmetric complex Gaussian distribution with a zero mean and an identity covariance matrix.
@@
where $\mathbf{y}=\left[y_{1},y_{2},\ldots,y_{M}\right]\in\mathbb{C}^{M\times 1}$ represents the signal received at the receiving array, $\mathbf{v}\in\mathbb{C}^{M\times 1}$ is the additive white Gaussian noise (AWGN), satisfying $\textbf{v}\sim\mathcal{CN}\left(\mathbf{0},\sigma^{2}_{v}\mathbf{I}\right)$, where $\sigma^{2}_{v}$ represents the average noise power.
@@ II-C Receiver
As shown in Fig. 1, $M$ antennas are arranged at the receiver in a uniform linear array with spacing $d_{r}$. Each antenna corresponds to an image class. In this letter, we aim to train the SIM-aided DNN to steer the wireless signals $\tilde{\mathbf{x}}$ towards the corresponding antenna. As a result, the image can be recognized by comparing the magnitude of the signal at each antenna. This process can be formulated as
@@
III SIM-aided DNN for Semantic Encoding
In this section, we first introduce the loss function used for evaluating the performance of the DNN. Then, the detailed training process of the SIM is explained.
@@ III-A Loss Function
The cross-entropy (CE) is used as the loss function to measure the difference between the power pattern across the receiving array and the expected probability distribution of recognizing the input image. Specifically, we use the softmax function to normalize the power at the receiving array, i.e., $\tilde{\mathbf{y}}=\text{softmax}(\mathbf{|y|}^{2})=\left[\tilde{y}_{1},\tilde{y}_{2},\ldots,\tilde{y}_{M}\right]\in\mathbb{R}^{M\times 1}$, where $\tilde{y}_{m}=\frac{e^{|{y_{m}}|^{2}}}{\sum_{i=1}^{M}e^{|{y_{i}}|^{2}}}$, $m\in\mathcal{M}$. Moreover, let $\mathbf{q}=\left[q_{1},q_{2},\ldots,q_{M}\right]\in\mathbb{R}^{M\times 1}$ denote the probability distribution, which is defined by
@@
III-B Training of DNN
In this subsection, we aim to train the transmission coefficients, including the amplitude vectors $\{\mathbf{a}^{1},\mathbf{a}^{2},\ldots,\mathbf{a}^{L}\}$ and phase vectors $\{\bm{\phi}^{1},\bm{\phi}^{2},\ldots,\bm{\phi}^{L}\}$, of the SIM for minimizing the total loss function over all image samples in the dataset. The training process mainly includes six steps:
@@ III-B Training of DNN
    Extract $B$ samples from the image dataset to create a mini-batch, and preprocess each image to match the size of the source encoding layer.
@@
After applying the mini-batch gradient descent method in several iterations, the loss function value converges. The detailed steps for training the SIM-aided DNN to enable image recognition tasks in the wave domain are given in Algorithm 1, where $\nabla$ in step 7 represents the gradient operator. Additionally, $\alpha>0$ in step $7$ is the learning rate, while $\mathbf{m}_{t}^{\{\mathbf{a}_{t}^{l},\bm{\phi}_{t}^{l}\}}$ and $\mathbf{v}_{t}^{\{\mathbf{a}_{t}^{l},\bm{\phi}_{t}^{l}\}}$ are respectively the first-order and second-order moment vectors. Additionally, $\beta_{1}$ and $\beta_{2}$ represent the decay coefficients, while $\epsilon$ is a small positive constant used for enhancing numerical stability. Appropriately adjusting these hyperparameters can strike tradeoffs between convergence speed and image recognition accuracy.
Algorithm 1 Training of the SIM-based DNN for enabling image recognition task-oriented semantic communications.
@@
9:  $\{\mathbf{a}^{1},\mathbf{a}^{2},\ldots,\mathbf{a}^{L}\}$, $\{\bm{\phi}^{1},\bm{\phi}^{2},\ldots,\bm{\phi}^{L}\}$.

### v2

@@ Abstract
Semantic communication (SemCom) leveraging advanced deep learning (DL) technologies enhances the efficiency and reliability of information transmission. Emerging stacked intelligent metasurface (SIM) with an electromagnetic neural network (EMNN) architecture enables complex computations at the speed of light. In this letter, we introduce an innovative SIM-aided SemCom system for image recognition tasks, where a SIM is positioned in front of the transmitting antenna. In contrast to conventional communication systems that transmit modulated signals carrying the image information or compressed semantic information, the carrier EM wave is directly transmitted from the source. The input layer of the SIM performs source encoding, while the remaining multi-layer architecture constitutes an EMNN for semantic encoding, transforming signals into a unique beam towards a receiving antenna corresponding to the image class. Remarkably, both the source and semantic encoding occur naturally as the EM waves propagate through the SIM. At the receiver, the image is recognized by probing the received signal magnitude across the receiving array. To this end, we utilize an efficient mini-batch gradient descent algorithm to train the transmission coefficients of SIM’s meta-atoms to learn the semantic representation of the image. Extensive numerical results verify the effectiveness of utilizing the SIM-based EMNN for image recognition task-oriented SemComs, achieving more than 90% recognition accuracy.
@@
Semantic communication (SemCom), stacked intelligent metasurface (SIM), image recognition, wave-domain computation.
@@ I Introduction
Conventional content-centric communication paradigms fail to satisfy the stringent quality-of-service requirements of next-generation wireless networks, particularly in terms of ultra-reliability, low latency, and high data rates [1]. As a result, there is a growing interest in intent-driven networks, which triggers the recent research surge in semantic communications (SemComs) [2, 3]. Relying on advanced artificial intelligence technologies and powerful computing capability, SemCom enables learning the semantic representations of various types of sources. For instance, the authors of [4] proposed a novel text SemCom system, DeepSC, to recover the meaning of sentences at the receiver with the minimum semantic errors. Based on this, several variants of DeepSC and advanced deep learning (DL)-driven SemCom systems have been developed [5, 6]. Nevertheless, the joint optimization of the semantic and channel encoders in these SemCom systems is typically time-consuming, which undoubtedly increases the training costs and deployment challenges for edges [6]. Additionally, the inference process would suffer from high computational complexity as the number of nodes increases.
@@
Recently, stacked intelligent metasurface (SIM) emerged as a promising technique for realizing fast and efficient wave-domain computation. Specifically, a SIM is a closed vacuum container with multiple closely stacked metasurface layers, where each metasurface consists of numerous low-cost meta-atoms whose electromagnetic (EM) response can be flexibly tuned by an intelligent controller, e.g., a field programmable gate array (FPGA) [7, 8]. As a result, SIM forms a programmable electromagnetic neural network (EMNN) in the EM domain, with signal processing implemented at the speed of light [9, 10, 11, 12]. Recently, researchers have utilized SIMs to perform various computing tasks in wireless communication and sensing scenarios, such as multiuser precoding [13, 14, 15, 16, 17] and direction-of-arrival (DOA) estimation [18, 19]. Nevertheless, the potential of utilizing SIM for SemComs has not hitherto been investigated.
@@
Against this background, we examine an image recognition task-oriented SemCom system leveraging a SIM. Compared to conventional SemCom systems sequentially encoding and transmitting semantic information, the wave-based computing paradigm enables automatic and efficient source and semantic encoding as the EM waves propagate through the SIM, greatly reducing energy consumption and processing latency. Specifically, in the proposed system, a SIM is positioned in front of the transmitting antenna, which only transmits the carrier EM waves. The input layer of the SIM performs image source encoding, while the remaining multi-layer architecture constitutes an EMNN for image semantic encoding, transforming the signals passing through the input layer of the SIM into a unique beam towards an antenna corresponding to the image class. As a result, the receiver recognizes the image by probing the signal magnitude across the receiving array. To this end, we utilize an efficient mini-batch gradient descent algorithm to train the transmission coefficients of meta-atoms in the SIM. Finally, extensive simulation results verify the effectiveness of utilizing a SIM to realize the task-oriented SemComs.
@@
Figure 1: A SIM-aided image recognition task-oriented SemCom system, where we consider $M$ image classes.
@@ II System Model
In this section, we introduce the SIM-aided SemCom system for image recognition. As shown in Fig. 1, we consider a system consisting of a single-antenna transmitter and a receiver equipped with $M$ antennas, which correspond to the number of image classes. Let =$\mathcal{M}=\left\{1,2,\ldots,M\right\}$ denote the set of indexes of antennas. More specifically, the transmitting antenna emits the unmodulated radio wave, and the SIM performs source and semantic coding of the image to direct the transmissive EM waves towards the corresponding antenna at the receiver.
@@ II-A1 Transmitting Antenna
In contrast to conventional communication systems, the transmitting antenna only radiates the carrier EM waves. Let $\sqrt{p_{t}}s$ denote the transmit signal, where $p_{t}$ represents the transmit power, and $s$ is the normalized EM signal [^1] Note that the receiver recognizes the image class by probing the amplitude of the EM signal, which means that the phase of $s$ has no impact on the recognition performance., satisfying $|s|^{2}=1$.
@@ II-A2 SIM
$w^{l}_{n,n^{*}}$ $=f_{w}\left(d_{s},d_{n,n^{*}}^{l},S_{a},\lambda\right)$
$=\frac{d_{s}S_{a}}{\left({d_{n,n^{*}}^{l}}\right)^{2}}\left(\frac{1}{2\pi d_{n,n^{*}}^{l}}-j\frac{1}{\lambda}\right)e^{j2\pi d_{n,n^{*}}^{l}},$ (1)
for $l\in\mathcal{L}/\{0\}$, $n^{*},n\in\mathcal{N}$, where $S_{a}$ is the area of each meta-atom, $d^{l}_{n,n^{*}}$ denotes the corresponding propagation distance between the two meta-atoms, $\lambda$ denotes the radio wavelength. Moreover, since the SIM is closely placed in front of the transmitter, the propagation channel from the transmitting antenna to the input layer of the SIM is characterized by $\mathbf{w}^{0}=f_{w}\left(d_{t,s},d^{0}_{n},S_{a},\lambda\right)\in\mathbb{C}^{N}$ [10], where $d_{t,s}$ and $d^{0}_{n}$ denote the horizontal distance from the transmitting antenna to the input layer of the SIM and the distance from the transmitter to the $n$-th meta-atom on the input layer, respectively.
@@
As illustrated in Fig. 1, the input layer of the SIM serves as a digital-to-analog converter to transform the pixel matrix of the source image into the transmission coefficients of the meta-atoms on the input layer [^2] For the sake of brevity, we consider only greyscale images in this letter. For color images, each pixel matrix of three channels can be preprocessed for fully leveraging the amplitude and phase components of the transmission coefficients of meta-atoms on the input layer.. Specifically, the image is first preprocessed to match the size of the input layer $\left(N_{\mathrm{row}}\times N_{\mathrm{col}}\right)$ by employing appropriate extraction or interpolation techniques [22], and then the transmission coefficients of the source encoding layer $\left(z^{0}_{1},z^{0}_{2},\ldots,z^{0}_{N}\right)$ are configured to the corresponding value of the pixel matrix. The amplitude and phase of the EM waves passing through the source encoding layer are determined by the product of the incident signal on the source coding layer, i.e., $\sqrt{p_{t}}s\mathbf{w}^{0}$ and the EM response imposed by the input layer, i.e., $\mathbf{Z}^{0}$. Therefore, the source encoded signal $\mathbf{x}\in\mathbb{C}^{N}$ radiated from the input metasurface layer can be expressed as
@@
The rest layers of the SIM construct an EMNN, which serves as a semantic encoder, aiming to transform $\mathbf{x}$ to a unique beam towards the receiving antenna corresponding to its image class. According to the Huygens–Fresnel principle [20], the EM wave passing through each meta-atom on the source encoding layer acts as a secondary source to illuminate all the meta-atoms on the first metasurface layer, and this process is repeated as the waves propagate through all metasurface layers [21]. Let $\mathbf{B}=\mathbf{Z}^{L}\mathbf{W}^{L}\ldots\mathbf{Z}^{2}\mathbf{W}^{2}\mathbf{Z}^{1}\mathbf{W}^{1}\in\mathbb{C}^{N\times N}$ denote the EM response of the semantic encoding layers. Hence, the semantic encoded signal $\tilde{\mathbf{x}}\in\mathbb{C}^{N}$ can be represented as
$\tilde{\mathbf{x}}=\mathbf{B}\mathbf{x}=\sqrt{p_{t}}s\mathbf{Z}^{L}\mathbf{W}^{L}\ldots\mathbf{Z}^{2}\mathbf{W}^{2}\mathbf{Z}^{1}\mathbf{W}^{1}\mathbf{Z}^{0}\mathbf{w}^{0}.$ (3)
@@ II-B Channel Model
where $p$ is the path loss, $K$ is the Rician factor, $\mathbf{H}_{\mathrm{LOS}}$ is the line-of-sight (LOS) component, characterizing the direct path between SIM and receiver, which is defined by ${[\mathbf{H}_{\mathrm{LOS}}]}_{m,n}=e^{-j\frac{2\pi}{\lambda}d_{m,n}}$, where $d_{m,n}$ denotes the distance from the $n$-th meta-atom on the output layer of the SIM to the $m$-th antenna at the receiving array. $\mathbf{H}_{\mathrm{NLOS}}$ is the non-LOS (NLOS) component, following the circularly symmetric complex Gaussian distribution with a zero mean and an identity covariance matrix. In this paper, we focus on the optimization of SIM to enable semantic encoding by assuming that the wireless channel coefficients have been estimated using advanced channel estimation approaches [11].
@@
where $\mathbf{y}=\left[y_{1},y_{2},\ldots,y_{M}\right]\in\mathbb{C}^{M}$ represents the signal received at the receiving array, $\mathbf{v}\in\mathbb{C}^{M}$ is the additive white Gaussian noise (AWGN), satisfying $\textbf{v}\sim\mathcal{CN}\left(\mathbf{0},\sigma^{2}_{v}\mathbf{I_{M}}\right)$, where $\mathbf{I_{M}}\in\mathbb{C}^{M\times M}$ denotes the identity matrix and $\sigma^{2}_{v}$ represents the average noise power.
@@ II-C Receiver
As shown in Fig. 1, $M$ antennas are arranged at the receiver in a uniform linear array with spacing $d_{r}$. Each antenna corresponds to an image class. In this letter, we aim to train the SIM-aided EMNN to steer the wireless signals $\tilde{\mathbf{x}}$ towards the corresponding antenna. As a result, the image can be recognized by comparing the magnitude of the signal at each antenna. This process can be formulated as
@@
III SIM-aided EMNN for Semantic Encoding
In this section, we first introduce the loss function used for evaluating the performance of the EMNN. Then, the detailed training process of the SIM is explained.
@@ III-A Loss Function
The cross-entropy (CE) is used as the loss function to measure the difference between the power pattern across the receiving array and the expected probability distribution of recognizing the input image. Specifically, we use the softmax function to normalize the power at the receiving array, i.e., $\tilde{\mathbf{y}}=\text{softmax}(\mathbf{|y|}^{2})=\left[\tilde{y}_{1},\tilde{y}_{2},\ldots,\tilde{y}_{M}\right]\in\mathbb{R}^{M}$, where $\tilde{y}_{m}=\frac{e^{|{y_{m}}|^{2}}}{\sum_{i=1}^{M}e^{|{y_{i}}|^{2}}}$, $m\in\mathcal{M}$. Moreover, let $\mathbf{q}=\left[q_{1},q_{2},\ldots,q_{M}\right]\in\mathbb{R}^{M}$ denote the probability distribution, which is defined by
@@
III-B Training of EMNN
In this subsection, we aim to train the transmission coefficients, including the amplitude vectors $\{\mathbf{a}^{1},\mathbf{a}^{2},\ldots,\mathbf{a}^{L}\}$ and phase vectors $\{\bm{\phi}^{1},\bm{\phi}^{2},\ldots,\bm{\phi}^{L}\}$, of the SIM for minimizing the total loss function of all image samples in the dataset [^3] For discrete phase shifts, one can optimize the SIM-aided EMNN using an alternating optimization approach [23].. The training process mainly includes six steps:
@@ III-B Training of EMNN
    Extract $B$ samples from the image dataset to create a mini-batch and preprocess each image to match the size of the source encoding layer.
@@
After applying the mini-batch gradient descent method in several iterations, the loss function value converges. The detailed steps for training the SIM-aided EMNN to enable image recognition tasks in the wave domain are given in Algorithm 1, where $\nabla$ in step 7 represents the gradient operator. Additionally, $\alpha>0$ in step $7$ is the learning rate, while $\mathbf{m}_{t}^{\{\mathbf{a}_{t}^{l},\bm{\phi}_{t}^{l}\}}$ and $\mathbf{v}_{t}^{\{\mathbf{a}_{t}^{l},\bm{\phi}_{t}^{l}\}}$ are respectively the first-order and second-order moment vectors. Additionally, $\beta_{1}$ and $\beta_{2}$ represent the decay coefficients, while $\epsilon$ is a small positive constant used for enhancing numerical stability. Appropriately adjusting these hyperparameters can strike tradeoffs between convergence speed and image recognition accuracy.
Algorithm 1 Training of the SIM-based EMNN for enabling image recognition task-oriented SemComs.
@@
9:  $\{\mathbf{a}^{1},\mathbf{a}^{2},\ldots,\mathbf{a}^{L}\}$, and $\{\bm{\phi}^{1},\bm{\phi}^{2},\ldots,\bm{\phi}^{L}\}$.